OpenStudy (anonymous):

PLZ HELPPPP SHOW WORK!!! 1. a^+5a+4/a^3 divided by a^2+3a+2/a^2-2a 2. 12x^2+9x-7 divided by 3x

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@reaper534 told me you could help me

OpenStudy (ranga):

In 1) a is raised to what?

OpenStudy (anonymous):

i have no clue

OpenStudy (ranga):

2) \[ \frac{12x^2 + 9x - 7}{3x} = \frac{12x^2}{3x} + \frac{9x}{3x} - \frac{7}{3x} = 4x + 3 - \frac{7}{3x}\]

OpenStudy (ranga):

In 1) go through the problem again. Something seems to be missing after a^ (the very first occurence).

OpenStudy (anonymous):

OS if finally working and ranga ill right out the question

OpenStudy (anonymous):

a^2

OpenStudy (anonymous):